# GEO: Generative Engine Optimization

### v2

@@ Abstract.
The advent of large language models (LLMs) has ushered in a new paradigm of search engines that use generative models to gather and summarize information to answer user queries. This emerging technology, which we formalize under the unified framework of generative engines (GEs), can generate accurate and personalized responses, rapidly replacing traditional search engines like Google and Bing. Generative Engines typically satisfy queries by synthesizing information from multiple sources and summarizing them using LLMs. While this shift significantly improves user utility and generative search engine traffic, it poses a huge challenge for the third stakeholder – website and content creators. Given the black-box and fast-moving nature of generative engines, content creators have little to no control over when and how their content is displayed. With generative engines here to stay, we must ensure the creator economy is not disadvantaged. To address this, we introduce Generative Engine Optimization (GEO), the first novel paradigm to aid content creators in improving their content visibility in GE responses through a flexible black-box optimization framework for optimizing and defining visibility metrics. We facilitate systematic evaluation by introducing GEO-bench, a large-scale benchmark of diverse user queries across multiple domains, along with relevant web sources to answer these queries. Through rigorous evaluation, we demonstrate that GEO can boost visibility by up to 40% in GE responses. Moreover, we show the efficacy of these strategies varies across domains, underscoring the need for domain-specific optimization methods. Our work opens a new frontier in information discovery systems, with profound implications for both developers of GEs and content creators.111Code and Data available at https://generative-engines.com/GEO/
@@ 1. Introduction
The invention of traditional search engines three decades ago revolutionized information access and dissemination globally. While they were powerful and ushered in a host of applications like academic research and e-commerce, they were limited to providing a list of relevant websites for user queries. However, the recent success of large language models (LLMs) has paved the way for better systems like BingChat, Google’s SGE, and perplexity.ai that combine conventional search engines with generative models. We dub these systems generative engines (GE) because they search for information and generate multi-modal responses by using multiple sources. Technically, generative engines (Figure 2) retrieve relevant documents from a database (like the internet) and use large neural models to generate a response grounded on the sources, ensuring attribution and a way for the user to verify the information.
The usefulness of generative engines for developers and users is evident – users access information faster and more accurately, while developers craft precise and personalized responses, improving user satisfaction and revenue. However, generative engines disadvantage the third stakeholder – website and content creators. Generative Engines, in contrast to traditional search engines, remove the need to navigate to websites by directly providing a precise and comprehensive response, potentially reducing organic traffic to websites and impacting their visibility. With millions of small businesses and individuals relying on online traffic and visibility for their livelihood, generative engines will significantly disrupt the creator economy. Further, the black-box and proprietary nature of generative engines makes it difficult for content creators to control and understand how their content is ingested and portrayed.
@@
To facilitate faithful and extensive evaluation of GEO methods, we propose GEO-bench, a benchmark consisting of 10000 queries from diverse domains and sources, adapted for generative engines. Through systematic evaluation, we demonstrate that our proposed Generative Engine Optimization methods can boost visibility by up to 40% on diverse queries, providing beneficial strategies for content creators. Among other things, we find that including citations, quotations from relevant sources, and statistics can significantly boost source visibility, with an increase of over 40% across various queries. We also demonstrate the efficacy of Generative Engine Optimization on Perplexity.ai, a state-of-the-art real-world generative engine and demonstrate visibility improvements up to 37%.
@@ 2.1. Formulation of Generative Engines
Generative Engines comprise two crucial components: a.) A set of generative models , each serving a specific purpose like query reformulation or summarization, and b.) A search engine that returns a set of sources given a query . We present a representative workflow in Figure 2, which at the time of writing, closely resembles the design of BingChat. This workflow breaks down the input query into a set of simpler queries that are easier to consume for the search engine. Given a query, a query re-formulating generative model, , generates a set of queries , which are then passed to the search engine to retrieve a set of ranked sources . The sets of sources are passed to a summarizing model , which generates a summary for each source in , resulting in the summary set (). The summary set is passed to a response-generating model , which generates a cumulative response backed by sources . In this work, we focus on single-turn Generative Engines, but the formulation can be extended to multi-turn Conversational Generative Engines (Appendix A).
@@ 2.2. Generative Engine Optimization
The advent of search engines led to search engine optimization (SEO), a process to help website creators optimize their content to improve search engine rankings. Higher rankings correlate with increased visibility and website traffic. However, traditional SEO methods are not directly applicable to Generative Engines. This is because, unlike traditional search engines, the generative model in generative engines is not limited to keyword matching, and the use of language models in ingesting source documents and response generation results in a more nuanced understanding of text documents and user query. With generative engines rapidly emerging as the primary information delivery paradigm and SEO not directly applicable, new techniques are needed. To this end, we propose Generative Engine Optimization, a new paradigm where content creators aim to increase their visibility (or impression) in generative engine responses. We define the visibility of a website (also referred as citation) in a cited response by the function , which the website creator wants to maximize. From the generative engine’s perspective, the goal is to maximize the visibility of citations most relevant to the user query, i.e., maximize ), where measures the relevance of citation to the query in the context of response and is determined by the exact algorithmic design of generative engine and is a black-box function to end-users. Further, both the functions and are subjective and not well-defined yet for generative engines, and we define them next.
@@ 2.2.1. Impressions for Generative Engines
To address this, we propose several impression metrics. The “Word Count” metric is the normalized word count of sentences related to a citation. Mathematically, this is defined as:
Here is the set of sentences citing , is the set of sentences in the response, and is the number of words in sentence . In cases where a sentence is cited by multiple sources, we share the word count equally with all the citations. Intuitively, a higher word count correlates with the source playing a more important part in the answer, and thus the user gets higher exposure to that source. However, since “Word Count” is not impacted by the ranking of the citations (whether it appears first, for example), we propose a position-adjusted count that reduces the weight by an exponentially decaying function of the citation position:
Intuitively, sentences that appear first in the response are more likely to be read, and the exponent term in definition gives higher weightage to such citations. Thus, a website cited at the top may have a higher impression despite having a lower word count than a website cited in the middle or end of the response. While the above impression metrics are objective and well-grounded, they ignore the subjective aspects of the impact of citations on the user’s attention. To address this, we propose the ”Subjective Impression” metric, which incorporates facets such as the relevance of the cited material to the user query, influence of the citation, uniqueness of the material presented by a citation, subjective position, subjective count, probability of clicking the citation, and diversity in the material presented. We use G-Eval (Liu et al., 2023a), the current state-of-the-art for evaluation with LLMs, to measure each of these sub-metrics.
@@ 2.2.2. Generative Engine Optimization methods for website
1: Authoritative: Modifies text style of the source content to be more persuasive and authoritative, 2. Statistics Addition: Modifies content to include quantitative statistics instead of qualitative discussion, wherever possible, 3. Keyword Stuffing: Modifies content to include more keywords from the query, as expected in classical SEO optimization. 4. Cite Sources & 5. Quotation Addition: Adds relevant citations and quotations from credible sources, 6.) 6. Easy-to-Understand: Simplifies the language of website, while 7. Fluency Optimization improves the fluency of website text. 8. Unique Words & 9. Technical Terms: involves adding unique and technical terms respectively wherever possible,
These methods cover diverse general strategies that website owners can implement quickly and use regardless of the website content. Further, except for methods 3, 4, and 5, the remaining methods enhance the presentation of existing content to increase its persuasiveness or appeal to the generative engine, without requiring extra content. On the other hand, methods 3,4, and 5 may require some form of additional content. To analyze the performance gain of our methods, for each input user query, we randomly select one source website to be optimized and apply each of the GEO methods separately on the same source. We refer readers to Appendix B.4 for exact prompts and more details on GEO methods.
@@ 3.1. Evaluated Generative Engine
In accordance with previous works (Liu et al., 2023b), we use a 2-step setup for Generative Engine design. The first step involves fetching relevant sources for input query, followed by a second step where an LLM generates a response based on the fetched sources. Due to context length limitations and quadratic scaling cost based on the context size of transformer models, only the top 5 sources are fetched from the Google search engine for every query. The setup closely mimics the workflow used in previous works and the general design adopted by commercial GEs such as you.com, perplexity.ai. The answer is then generated by gpt3.5-turbo model using the prompt same as prior work (Liu et al., 2023b). We refer readers to Appendix B for more details.
@@ 3.2. Benchmark : GEO-bench
Since there is currently no publicly available dataset containing Generative Engine related queries, we curate GEO-bench, a benchmark consisting of 10K queries from multiple sources, repurposed for generative engines, along with synthetically generated queries. The benchmark includes queries from nine different sources; each further categorized based on their target domain, difficulty, query intent, and other dimensions.
@@
1. MS Macro, 2. ORCAS-1, and 3. Natural Questions: (Kwiatkowski et al., 2019; Alexander et al., 2022; Craswell et al., 2021) These datasets contain real anonymized user queries from Bing and Google Search Engines. These three collectively represent the common set of datasets that are used in search engine related research. However, Generative Engines will be posed with far more difficult and specific queries with the intent of synthesizing answers from multiple sources instead of searching for them. To this end, we repurpose several other publicly available datasets: 4. AllSouls: This dataset contains essay questions from ”All Souls College, Oxford University”. The queries in this dataset require Generative Engines to perform appropriate reasoning to aggregate information from multiple sources. 5. LIMA: (Zhou et al., 2023) contains challenging questions requiring Generative Engines to not only aggregate information but also perform suitable reasoning to answer the question (eg: writing a short poem, python code.). 6. Davinci-Debtate (Liu et al., 2023b) contains debate questions generated for testing Generative Engines. 7. Perplexity.ai Discover222https://www.perplexity.ai/discover: These queries are sourced from Perplexity.ai’s Discover section, which is an updated list of trending queries on the platform. 8. ELI-5333https://huggingface.co/datasets/eli5_category: This dataset contains questions from the ELI5 subreddit, where users ask complex questions and expect answers in simple, layman’s terms. 9. GPT-4 Generated Queries: To supplement diversity in query distribution, we prompt GPT-4 to generate queries ranging from various domains (eg, science, history) and based on query intent (eg, navigational, transactional) and based on difficulty and scope of generated response (eg, open-ended, fact-based).
@@
3.3. GEO Methods
3.4. Evaluation Metrics
Furthermore, all impression metrics are normalized by multiplying them with a constant factor so that the sum of the impressions of all citations in a response equals 1. In our discussion, we compare methods by calculating the relative improvement in impression. For an initial generated response from sources , and a modified response , the relative improvement in impression for each source is measured as:
@@ 4. Results
We evaluate various Generative Engine Optimization methods designed to optimize website content for better visibility in Generative Engine responses, compared against a baseline with no optimization. Our evaluation used GEO-bench, a diverse benchmark of user queries from multiple domains and settings. Performance was measured using two metrics: Position-Adjusted Word Count and Subjective Impression. The former consider word count and citation position in the GE’s response, while the latter computes multiple subjective factors, giving an overall impression score.
@@ 5.1. Domain-Specific Generative Engine Optimizations
Similarly, the addition of citations through Cite Sources is particularly beneficial for factual questions, likely because citations provide a source of verification for the facts presented, thereby enhancing the credibility of the response. The effectiveness of different GEO methods varies across domains. For example, as shown in row 5 of Table 3, domains such as ‘Law & Government’ and question types like ‘Opinion’ benefit significantly from the addition of relevant statistics in the website content, as implemented by Statistics Addition. This suggests that data-driven evidence can enhance the visibility of a website in particular contexts. The method Quotation Addition is most effective in the ‘People & Society’, ‘Explanation’, and ‘History’ domains. This could be because these domains often involve personal narratives or historical events, where direct quotes can add authenticity and depth to the content. Overall, our analysis suggests that website owners should strive towards making domain-specific targeted adjustments to their websites for higher visibility.
@@ 5.4. Qualitative Analysis
| Method | GEO Optimization | Relative Improvement |
| Query: What is the secret of Swiss chocolate | ||
| Cite Sources | 132.4% | |
| Query: Should robots replace humans in the workforce? | ||
| Source: Not here, and not now — until recently. The big difference is that robots have come not to destroy our lives, but to disrupt our work, with a staggering 70% increase in robotic involvement in last decade. | ||
| Statistics Addition | 65.5% | |
| Query: Did the jacksonville jaguars ever make it to the superbowl? | ||
| Authoritative | 89.1% |
@@ 7. Related Work
Previous works have used several techniques for answer generation backed by sources. Nakano et al. (2021) trained GPT-3 to navigate web environments to generate source-backed answers. Similarly, other methods (Shuster et al., 2022; Thoppilan et al., 2022; Menick et al., 2022) fetch sources via search engines for answer generation. Our work unifies these approaches and provide a common benchmark for improving these systems in the future. In a recent working draft, Kumar and Lakkaraju (Kumar and Lakkaraju, 2024) showed that strategic text sequences can manipulate LLM recommendations to enhance product visibility in generative engines. While their approach focuses on increasing product visibility through adversarial text, our method introduces non-adversarial strategies to optimize any website content for improved visibility in generative engine search results.

### v3

@@ Abstract.
The advent of large language models (LLMs) has ushered in a new paradigm of search engines that use generative models to gather and summarize information to answer user queries. This emerging technology, which we formalize under the unified framework of generative engines (GEs), can generate accurate and personalized responses, rapidly replacing traditional search engines like Google and Bing. Generative Engines typically satisfy queries by synthesizing information from multiple sources and summarizing them using LLMs. While this shift significantly improves user utility and generative search engine traffic, it poses a huge challenge for the third stakeholder – website and content creators. Given the black-box and fast-moving nature of generative engines, content creators have little to no control over when and how their content is displayed. With generative engines here to stay, we must ensure the creator economy is not disadvantaged. To address this, we introduce Generative Engine Optimization (GEO), the first novel paradigm to aid content creators in improving their content visibility in generative engine responses through a flexible black-box optimization framework for optimizing and defining visibility metrics. We facilitate systematic evaluation by introducing GEO-bench, a large-scale benchmark of diverse user queries across multiple domains, along with relevant web sources to answer these queries. Through rigorous evaluation, we demonstrate that GEO can boost visibility by up to 40% in generative engine responses. Moreover, we show the efficacy of these strategies varies across domains, underscoring the need for domain-specific optimization methods. Our work opens a new frontier in information discovery systems, with profound implications for both developers of generative engines and content creators.111Code and Data available at https://generative-engines.com/GEO/
@@ 1. Introduction
The invention of traditional search engines three decades ago revolutionized information access and dissemination globally (Brin and Page, 1998). While they were powerful and ushered in a host of applications like academic research and e-commerce, they were limited to providing a list of relevant websites for user queries. However, the recent success of large language models (Brown et al., 2020; OpenAI et al., 2024) has paved the way for better systems like BingChat, Google’s SGE, and perplexity.ai that combine conventional search engines with generative models. We dub these systems generative engines (GE) because they search for information and generate multi-modal responses by using multiple sources. Technically, generative engines (Figure 2) retrieve relevant documents from a database (like the internet) and use large neural models to generate a response grounded on the sources, ensuring attribution and a way for the user to verify the information.
The usefulness of generative engines for developers and users is evident – users access information faster and more accurately, while developers craft precise and personalized responses, improving user satisfaction and revenue. However, generative engines disadvantage the third stakeholder – website and content creators. Generative Engines, in contrast to traditional search engines, remove the need to navigate to websites by directly providing a precise and comprehensive response, potentially reducing organic traffic to websites and impacting their visibility (Maayan, 2023). With millions of small businesses and individuals relying on online traffic and visibility for their livelihood, generative engines will significantly disrupt the creator economy. Further, the black-box and proprietary nature of generative engines makes it difficult for content creators to control and understand how their content is ingested and portrayed.
@@
To facilitate faithful and extensive evaluation of GEO methods, we propose GEO-bench, a benchmark consisting of 10000 queries from diverse domains and sources, adapted for generative engines. Through systematic evaluation, we demonstrate that our proposed Generative Engine Optimization methods can boost visibility by up to 40% on diverse queries, providing beneficial strategies for content creators. Among other things, we find that including citations, quotations from relevant sources, and statistics can significantly boost source visibility, with an increase of over 40% across various queries. We also demonstrate the efficacy of Generative Engine Optimization on Perplexity.ai, a real-world generative engine and demonstrate visibility improvements up to 37%.
@@ 2.1. Formulation of Generative Engines
Generative Engines comprise two crucial components: a.) A set of generative models , each serving a specific purpose like query reformulation or summarization, and b.) A search engine that returns a set of sources given a query . We present a representative workflow in Figure 2, which, at the time of writing, closely resembles the design of BingChat. This workflow breaks down the input query into a set of simpler queries that are easier to consume for the search engine. Given a query, a query re-formulating generative model, , generates a set of queries , which are then passed to the search engine to retrieve a set of ranked sources . The sets of sources are passed to a summarizing model , which generates a summary for each source in , resulting in the summary set (). The summary set is passed to a response-generating model , which generates a cumulative response backed by sources . In this work, we focus on single-turn Generative Engines, but the formulation can be extended to multi-turn Conversational Generative Engines (Appendix A).
@@ 2.2. Generative Engine Optimization
The advent of search engines led to search engine optimization (SEO), a process to help website creators optimize their content to improve search engine rankings. Higher rankings correlate with increased visibility and website traffic. However, traditional SEO methods are not directly applicable to Generative Engines. This is because, unlike traditional search engines, the generative model in generative engines is not limited to keyword matching, and the use of language models in ingesting source documents and response generation results in a more nuanced understanding of text documents and user query. With generative engines rapidly emerging as the primary information delivery paradigm and SEO is not directly applicable; new techniques are needed. To this end, we propose Generative Engine Optimization, a new paradigm where content creators aim to increase their visibility (or impression) in generative engine responses. We define the visibility of a website (also referred to as a citation) in a cited response by the function , which the website creator wants to maximize. From the generative engine’s perspective, the goal is to maximize the visibility of citations most relevant to the user query, i.e., maximize ), where measures the relevance of citation to the query in the context of response and is determined by the exact algorithmic design of generative engine and is a black-box function to end-users. Further, both the functions and are subjective and not well-defined yet for generative engines, and we define them next.
@@ 2.2.1. Impressions for Generative Engines
In response to this challenge, we propose a suite of impression metrics designed with three key principles in mind: 1.) The metrics should hold relevance for creators, 2.) They should be explainable, and 3.) They should be easily comprehensible by a broad spectrum of content creators. The first of these metrics, the “Word Count” metric, is the normalized word count of sentences related to a citation. Mathematically, this is defined as:
Here is the set of sentences citing , is the set of sentences in the response, and is the number of words in sentence . In cases where a sentence is cited by multiple sources, we share the word count equally with all the citations. Intuitively, a higher word count correlates with the source playing a more important part in the answer, and thus, the user gets higher exposure to that source. However, since “Word Count” is not impacted by the ranking of the citations (whether it appears first, for example), we propose a position-adjusted count that reduces the weight by an exponentially decaying function of the citation position:
Intuitively, sentences that appear first in the response are more likely to be read, and the exponent term in definition gives higher weightage to such citations. Thus, a website cited at the top may have a higher impression despite having a lower word count than a website cited in the middle or end of the response. Further, the choice of exponentially decaying function is motivated by several studies showing click-through rates follow a power-law as a function of ranking in search engines (Goodwin, 2011; Dean, 2023). While the above impression metrics are objective and well-grounded, they ignore the subjective aspects of the impact of citations on the user’s attention. To address this, we propose the ”Subjective Impression” metric, which incorporates facets such as the relevance of the cited material to the user query, influence of the citation, uniqueness of the material presented by a citation, subjective position, subjective count, probability of clicking the citation, and diversity in the material presented. We use G-Eval (Liu et al., 2023a), the current state-of-the-art for evaluation with LLMs, to measure each of these sub-metrics.
@@ 2.2.2. Generative Engine Optimization methods for website
1: Authoritative: Modifies text style of the source content to be more persuasive and authoritative, 2. Statistics Addition: Modifies content to include quantitative statistics instead of qualitative discussion, wherever possible, 3. Keyword Stuffing: Modifies content to include more keywords from the query, as expected in classical SEO optimization. 4. Cite Sources & 5. Quotation Addition: Adds relevant citations and quotations from credible sources respectively, 6.) 6. Easy-to-Understand: Simplifies the language of website, while 7. Fluency Optimization improves the fluency of website text. 8. Unique Words & 9. Technical Terms: involves adding unique and technical terms respectively wherever possible,
These methods cover diverse general strategies that website owners can implement quickly and use regardless of the website content. Further, except for methods 3, 4, and 5, the remaining methods enhance the presentation of existing content to increase its persuasiveness or appeal to the generative engine, without requiring extra content. On the other hand, methods 3,4 and 5 may require some form of additional content. To analyze the performance gain of our methods, for each input user query, we randomly select one source website to be optimized and apply each of the GEO methods separately on the same source. We refer readers to Appendix B.4 for more details on GEO methods.
@@ 3.1. Evaluated Generative Engine
In accordance with previous works (Liu et al., 2023b), we use a 2-step setup for Generative Engine design. The first step involves fetching relevant sources for input query, followed by a second step where an LLM generates a response based on the fetched sources. Similar to prevous works, we do not use summarization and provide the whole response for each source. Due to context length limitations and quadratic scaling cost based on the context size of transformer models, only the top 5 sources are fetched from the Google search engine for every query. The setup closely mimics the workflow used in previous works and the general design adopted by commercial GEs such as you.com and perplexity.ai. The answer is then generated by the gpt3.5-turbo model (OpenAI, 2022) using the same prompt as prior work (Liu et al., 2023b). We sample 5 different responses at temperature=0.7, to reduce statistical deviations.
@@ 3.2. Benchmark : GEO-bench
Since there is currently no publicly available dataset containing Generative Engine related queries, we curate GEO-bench, a benchmark consisting of 10K queries from multiple sources, repurposed for generative engines, along with synthetically generated queries. The benchmark includes queries from nine different sources, each further categorized based on their target domain, difficulty, query intent, and other dimensions.
@@
1. MS Macro, 2. ORCAS-1, and 3. Natural Questions: (Kwiatkowski et al., 2019; Alexander et al., 2022; Craswell et al., 2021) These datasets contain real anonymized user queries from Bing and Google Search Engines. These three collectively represent the common set of datasets that are used in search engine related research. However, Generative Engines will be posed with far more difficult and specific queries with the intent of synthesizing answers from multiple sources instead of searching for them. To this end, we repurpose several other publicly available datasets: 4. AllSouls: This dataset contains essay questions from ”All Souls College, Oxford University.” The queries in this dataset require Generative Engines to perform appropriate reasoning to aggregate information from multiple sources. 5. LIMA: (Zhou et al., 2023) contains challenging questions requiring Generative Engines to not only aggregate information but also perform suitable reasoning to answer the question (e.g., writing a short poem, python code.). 6. Davinci-Debtate (Liu et al., 2023b) contains debate questions generated for testing Generative Engines. 7. Perplexity.ai Discover222https://www.perplexity.ai/discover: These queries are sourced from Perplexity.ai’s Discover section, which is an updated list of trending queries on the platform. 8. ELI-5333https://huggingface.co/datasets/eli5_category: This dataset contains questions from the ELI5 subreddit, where users ask complex questions and expect answers in simple, layman’s terms. 9. GPT-4 Generated Queries: To supplement diversity in query distribution, we prompt GPT-4 (OpenAI et al., 2024) to generate queries ranging from various domains (e.g., science, history) and based on query intent (e.g., navigational, transactional) and based on difficulty and scope of generated response (e.g., open-ended, fact-based).
@@
3.3. GEO Methods
3.4. Evaluation Metrics
Furthermore, all impression metrics are normalized by multiplying them with a constant factor so that the sum of the impressions of all citations in a response equals 1. In our analysis, we compare methods by calculating the relative improvement in impression. For an initial generated response from sources , and a modified response , the relative improvement in impression for each source is measured as:
@@ 4. Results
We evaluate various Generative Engine Optimization methods designed to optimize website content for better visibility in Generative Engine responses, compared against a baseline with no optimization. Our evaluation used GEO-bench, a diverse benchmark of user queries from multiple domains and settings. Performance was measured using two metrics: Position-Adjusted Word Count and Subjective Impression. The former considers word count and citation position in the GE’s response, while the latter computes multiple subjective factors, giving an overall impression score.
@@ 5.1. Domain-Specific Generative Engine Optimizations
Similarly, the addition of citations through Cite Sources is particularly beneficial for factual questions, likely because citations provide a source of verification for the facts presented, thereby enhancing the credibility of the response. The effectiveness of different GEO methods varies across domains. For example, as shown in row 5 of Table 3, domains such as ‘Law & Government’ and question types like ‘Opinion’ benefit significantly from the addition of relevant statistics in the website content, as implemented by Statistics Addition. This suggests that data-driven evidence can enhance the visibility of a website in particular contexts. The method Quotation Addition is most effective in the ‘People & Society,’ ‘Explanation,’ and ‘History’ domains. This could be because these domains often involve personal narratives or historical events, where direct quotes can add authenticity and depth to the content. Overall, our analysis suggests that website owners should strive towards making domain-specific targeted adjustments to their websites for higher visibility.
| Method | GEO Optimization | Relative Improvement |
| Query: What is the secret of Swiss chocolate | ||
| Cite Sources | 132.4% | |
| Query: Should robots replace humans in the workforce? | ||
| Source: Not here, and not now — until recently. The big difference is that the robots have come not to destroy our lives, but to disrupt our work, with a staggering 70% increase in robotic involvement in the last decade. | ||
| Statistics Addition | 65.5% | |
| Query: Did the jacksonville jaguars ever make it to the superbowl? | ||
| Authoritative | 89.1% |
@@ 7. Related Work
Previous works have used several techniques for answer generation backed by sources. Nakano et al. (2021) trained GPT-3 to navigate web environments to generate source-backed answers. Similarly, other methods (Shuster et al., 2022; Thoppilan et al., 2022; Menick et al., 2022) fetch sources via search engines for answer generation. Our work unifies these approaches and provides a common benchmark for improving these systems in the future. In a recent working draft, Kumar and Lakkaraju (Kumar and Lakkaraju, 2024) showed that strategic text sequences can manipulate LLM recommendations to enhance product visibility in generative engines. While their approach focuses on increasing product visibility through adversarial text, our method introduces non-adversarial strategies to optimize any website content for improved visibility in generative engine search results.